OpenStudy (anonymous):

Use the ratio test and find the radius of convergence of the following series

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}\frac{ x ^{n} }{ n ^{\frac{ 3 }{ 2 }}3^{n} }\]

OpenStudy (anonymous):

You must find this limit: \[\lim_{n \rightarrow \infty}\frac{ a _{n+1}^{} }{ a _{n} }\]

OpenStudy (anonymous):

do you have troubles substituing?

OpenStudy (anonymous):

Yeah, I know, but I wasnt sure which terms I could cancel as n->infinity

OpenStudy (anonymous):

well I'll show you

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):